OpenStudy (anonymous):

Please Solve Integral (4-4cos4x)^5/2 dx

OpenStudy (anonymous):

Is the exponent 5/2, or is the exponent just 5 and the whole expression is divided by 2?

OpenStudy (anonymous):

int(4-4cos4x)^5/2dx =4^5/2int(1-cos4x)5/2dx =4^5/2int(2cos^2(2x))^5/2dx then use reduction formula for int cos^(n)x